OpenStudy (anonymous):

Solve using the addition method. x+y=12 2x-y=6

OpenStudy (amistre64):

is the additionmethod the same as the elimination method?

OpenStudy (anonymous):

Add the two to get 3x = 18 x = 18/3 = 6 Put in first equation 6 + y = 12 y = 12 - 6 = 6